an 80 ft flagpole has upper and lowe diameters of 4 and 16 in, respectively. find the cost of painting it at 10 cents per square foot.

Answer by mananth(16946)


vertical height of pole = 80 ft
diameters of 4 = 2 in = 1/6 ft
radius
and 16 in , 8 inch radius = 3/4 ft
Curved surface area = pi * (r1 +r2) h
= pi * (1/2)*(11/12) * 80
= 115.2 ft^2
multiply by 10 to get the cost